Transaction costs include

costs of negotiating contracts with other firms.

cost of enforcing contracts.

the existence of asset-specificity.

all of the above

1.25000 points   

QUESTION 8

Firms are organized to keep their costs as low as possible by

comparing internal operating cost and external transactions cost associated with outsourcing.

analyzing supply and demand conditions.

minimizing their use of borrowed funds.

pricing their products appropriately.

1.25000 points   

QUESTION 9

Consider the following graph, which shows two lines. The downward-sloping line represents the demand curve (the quantity demanded at various prices) and the upward-sloping line represents the supply curve (the quantity supplied at various prices).

Supply and Demand $10 $9 $8 은$7 $6 $5 $4 Quantity Supplied -Quettity Demanded Quantity 4 10 12

Which price corresponds to a quantity demanded of 10?

$ 8

$ 7

$ 6

$ 5

1.25000 points   

QUESTION 10

Consider the following graph, which shows two lines. The downward-sloping line represents the demand curve (the quantity demanded at various prices) and the upward-sloping line represents the supply curve (the quantity supplied at various prices).

_jn5v5Bhilsa--9YkaFgJIiyaAedTLZM8wMhl2P1

What is the difference between the quantity demanded and the quantity supplied when the price is $5?

4

6

8

10

1.25000 points   

QUESTION 11

Consider the following graph, which shows two lines. The downward-sloping line represents the demand curve (the quantity demanded at various prices) and the upward-sloping line represents the supply curve (the quantity supplied at various prices).

_jn5v5Bhilsa--9YkaFgJIiyaAedTLZM8wMhl2P1

At which price is the quantity demanded equal to the quantity supplied?

$ 8

$ 7

$ 6

$ 5

1.25000 points   

QUESTION 12

Consider the following table, which lists the quantity demanded and the quantity supplied at various prices.

Price

Quantity Demanded

Quantity Supplied

$5

120

20

10

90

40

15

60

60

20

30

80

25

0

100

At which price are the quantity demanded and quantity supplied the same?

$10

$15

$20

None of the above

1.25000 points   

QUESTION 13

Consider the following table, which lists the quantity demanded and the quantity supplied at various prices.

Price

Quantity Demanded

Quantity Supplied

$5

120

20

10

90

40

15

60

60

20

30

80

25

0

100

At which price is the quantity demanded more than the quantity supplied (a shortage)?

$10

$15

$20

None of the above

1.25000 points   

QUESTION 14

Consider the following table, which lists the quantity demanded and the quantity supplied at various prices.

Price

Quantity Demanded

Quantity Supplied

$5

120

20

10

90

40

15

60

60

20

30

80

25

0

100

At which price is the quantity supplied more than the quantity demanded (a surplus)??

$10

$15

$20

None of the above

1.25000 points   

QUESTION 15

Consider the following equations, which represent quantity demanded and quantity supplied in a market for coffee mugs.

Quantity Supplied (QS) = P + 2

Quantity Demanded (QD) = ̶ 2P + 20

How many mugs would be supplied and how many would be demanded at a price of $5?

2 supplied, 18 demanded

2 supplied, 20 demanded

7 supplied, 10 demanded

7 supplied, 18 demanded

1.25000 points   

QUESTION 16

Consider the following equations, which represent quantity demanded and quantity supplied in a market for coffee mugs.

Quantity Supplied (QS) = P + 2

Quantity Demanded (QD) = ̶ 2P + 20

How many mugs would be supplied and how many would be demanded at a price of $6?

6 supplied, 8 demanded

6 supplied, 18 demanded

8 supplied, 8 demanded

8 supplied, 18 demanded
Answer #1

Q7. Answer is Cost of negotiating the ccontracts with other firms.

Q8. Answer is Comparing internal operating cost and external transaction cost associated with outsourcing.

Q9. Answer is $5.

Q10. Answer is 6.

Explanation: At price of $5

Supply: 4

Demand:10

Difference between demand and supply: 10-4 = 6

Q11. Aanswer is $7.

Explanation: At price $7, both demandd and supply is 8 units, as reflecting from graph.

Q12. Answer is $15.

Q13. Answer is $10.

Explanation: At price at 10,

Demand: 90

Supply: 40

Shortage: Demand-Supply = 90 -40 = 50

Q14. Answer is $20

Explanation: At price $20:

Supply: 80 Demand: 30

Surplus: Supply - Demand = 80-30 = 50

Q15. Answer is 7 suplied and 10 demanded

Explanation:

QS = P+2 ,when Price =5, QS =7 Units

QD= -2P +20 , when Price =5, QD =10 units

Q16. Answer is 8 supplied and 8 demanded

Explanation: At price $6,

QS = P+2 when Price =6 , QS = 8 units

QD = -2P+20, when price =6, QD =8 units